# MixedGaussianAvatar: Realistically and Geometrically Accurate Head Avatar via Mixed 2D-3D Gaussian Splatting

### v2

@@ Abstract
Reconstructing high-fidelity 3D head avatars is crucial in various applications such as virtual reality. The pioneering methods reconstruct realistic head avatars with Neural Radiance Fields (NeRF), which have been limited by training and rendering speed. Recent methods based on 3D Gaussian Splatting (3DGS) significantly improve the efficiency of training and rendering. However, the surface inconsistency of 3DGS results in subpar geometric accuracy; later, 2DGS uses 2D surfels to enhance geometric accuracy at the expense of rendering fidelity. To leverage the benefits of both 2DGS and 3DGS, we propose a novel method named MixedGaussianAvatar for realistically and geometrically accurate head avatar reconstruction. Our main idea is to utilize 2D Gaussians to reconstruct the surface of the 3D head, ensuring geometric accuracy. We attach the 2D Gaussians to the triangular mesh of the FLAME model and connect additional 3D Gaussians to those 2D Gaussians where the rendering quality of 2DGS is inadequate, creating a mixed 2D-3D Gaussian representation. These 2D-3D Gaussians can then be animated using FLAME parameters. We further introduce a progressive training strategy that first trains the 2D Gaussians and then fine-tunes the mixed 2D-3D Gaussians. We demonstrate the superiority of MixedGaussianAvatar through comprehensive experiments. The code will be released at: https://github.com/ChenVoid/MGA/.
22footnotetext: Corresponding Author.
Introduction
@@
Reconstructing high-quality 3D head avatars from multiple 2D images captured from different viewpoints is essential for various applications (Wohlgenannt, Simons, and Stieglitz 2020; Chollet et al. 2009; Bai et al. 2024). The key high-fidelity aspects include the geometric and realistic accuracy of 3D head avatars and rendered images. With the advancement of deep learning, techniques based on Neural Radiance Fields (NeRF) (Mildenhall et al. 2021) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) have gained popularity, providing significant benefits for creating high-fidelity 3D head avatars.
@@
Figure 1: MixedGaussianAvatar uses a mixed 2D-3D Gaussian Splatting method to reconstruct a realistically and geometrically accurate 3D head avatar mesh.
@@
NeRF introduced the idea of radiance fields, utilizing neural networks to store 3D information, which allows for reconstructing high-quality static scenes. When it comes to 3D head avatar reconstruction, the proposed methods primarily focus on reconstructing dynamic head avatars (Gao et al. 2022; Zielonka, Bolkart, and Thies 2022). For example, Neural Head Avatars (NHA) (Grassal et al. 2022) process RGB video inputs that display various expressions and perspectives. NHA involves estimating and refining the low-dimensional shape, expression, and pose parameters of the FLAME head model from an RGB input frame, which allows for creating a neural head avatar that can be animated and rendered from novel viewpoints. However, the head avatars generated using NeRF-based methods have lower accuracy due to the implicit representation characteristics of NeRF, and the training and rendering processes are slow.
@@
Recently, 3DGS effectively modeled a static scene using 3D Gaussians and rendered it through a differentiable rasterizer, greatly speeding up the 3D reconstruction process from multi-view RGB images. Because of 3DGS’s explicit representation and efficient rasterization properties, it achieves high rendering quality and fast training and inference speeds. In 3D head avatar reconstruction, most methods attach 3D Gaussians to the FLAME triangular mesh to create dynamic facial expressions. For example, GaussianAvatars (Qian et al. 2024) employ a rigging approach to align the local positions, rotations, and scales of 3D Gaussians with the global parameters. FlashAvatar (Xiang et al. 2024) and Gaussian Head Avatar  (Xu et al. 2024) utilize neural networks to predict 3D Gaussian parameters. The advantage of these methods lies in their ability to dynamically represent 3D head avatars while producing high-quality images. However, due to the inherent multi-view inconsistency of 3DGS, it cannot accurately reconstruct geometric surfaces.
@@
2D Gaussian Splatting (Huang et al. 2024a) replaces 3D Gaussians with 2D Gaussians and employs a ray-splat intersection method for the splatting process, ensuring consistency across multiple views. The advantage lies in the high quality of reconstructed geometric surfaces, but it compromises the quality of rendered images. To summarize 3DGS and 2DGS, 1) 3DGS has achieved realistic results in novel view synthesis, but it struggles to capture accurate geometric structures. 2) 2DGS uses 2D surfels to accurately reconstruct geometric surfaces, but it cannot render realistic images as effectively as 3DGS.
@@
To harness the advantages of both 3DGS and 2DGS, we introduce an innovative method called MixedGaussianAvatar. This method uses a mixed 2D-3D Gaussians representation, combining the color rendering advantages of 3DGS with the geometric reconstruction strengths of 2DGS to achieve a realistically and geometrically accurate 3D head avatar reconstruction. As shown in Fig. 1, our primary goal is to use 2DGS to preserve the geometric accuracy of the 3D head surface. We attach 2D Gaussians to the triangular mesh of the FLAME model and then connect additional 3D Gaussians to the corresponding 2D Gaussians in areas where the rendering quality is insufficient. The mixed 2D-3D Gaussians can be driven using FLAME parameters to form a dynamic 3D representation. To train the mixed 2D-3D Gaussians, we further propose a progressive training strategy that first trains the 2D Gaussians and then finetunes the mixed 2D-3D Gaussians. Our contributions are summarized as follows:
@@
Related Work
Figure 2: Pipeline of MixedGaussianAvatar. We propose a differentiable mixed 2D-3D Gaussian Splatting method for reconstructing realistic and geometrically accurate head avatars from multi-view 2D images. This approach uses 2D Gaussians for geometric precision and 3D Gaussians to correct color errors, resulting in high-quality 3D head mesh and realistic mesh textures or RGB images. The method is integrated with the FLAME model, enabling dynamic effects and animation driven by parameters through local-to-global transformation and the progressive training strategy.
@@
The representation of dynamic scenes has attracted considerable attention. Previous methods, like NeRF (Mildenhall et al. 2021) and its variants (Müller et al. 2022; Chen et al. 2022; Wei et al. 2024), focus on modeling static scenes, storing them in learnable neural networks to achieve competitive novel view synthesis quality. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) and its variants (Lu et al. 2024) have introduced a much faster training and rendering pipeline. The methods mentioned above are only effective for modeling static scenes and have difficulty handling dynamic scenes. Based on NeRF and 3DGS, several approaches (Shao et al. 2023; Wu et al. 2024; Huang et al. 2024b) introduce 4D coordinates to effectively represent dynamic scenes. Humans are very sensitive to facial details, making it challenging for general methods to create high-fidelity head avatars. Recently, numerous efforts (Grassal et al. 2022; Hong et al. 2022; Chen et al. 2023) have been made to develop specialized techniques for representing 3D head avatars. AvatarMAV (Xu et al. 2023) and INSTA (Zielonka, Bolkart, and Thies 2023), both based on voxel representations, enable fast avatar rendering. PointAvatar (Zheng et al. 2023) is the first to employ a deformable point-based representation for creating high-fidelity avatars. Inspired by 3DGS and PointAvatar, methods like GaussianAvatars (Qian et al. 2024), Gaussian Head Avatar (Xu et al. 2024), and FlashAvatar (Xiang et al. 2024) attach 3D Gaussians to dynamic meshes or UV maps, achieving success in real-time 3D head avatar animation. Previous methods fail to consider geometric reconstruction, which makes them less appropriate for industrial applications. In contrast, MixedGaussianAvatar integrates the high-quality appearance of 3DGS with the consistent geometry of 2DGS.
@@
Compared to neural rendering, which primarily focuses on view synthesis, neural reconstruction emphasizes explicit surface extraction and textured mesh reconstruction. NeuS (Wang et al. 2021) is the first to propose a bias-free volume rendering method based on neural fields that leverages a signed distance function (SDF) to achieve high-quality surface reconstruction. UNISURF (Oechsle, Peng, and Geiger 2021) presents a unified framework that combines neural implicit surfaces and radiance fields to enable accurate 3D surface reconstruction from multi-view images without requiring object masks. VolSDF (Yariv et al. 2021) significantly improves geometry approximation and disentangles shape and appearance compared to previous neural rendering methods. With the development of 3D Gaussian Splatting, SuGaR (Guédon and Lepetit 2024) introduces a method for fast and precise 3D mesh reconstruction and high-quality mesh rendering by aligning 3D Gaussian splats with scene surfaces, enabling efficient and detailed mesh extraction within minutes. 2D Gaussian Splatting (2DGS) (Huang et al. 2024a) proposes a novel approach for geometrically accurate radiance field reconstruction by leveraging 2D Gaussian primitives, which significantly improves surface modeling and view-consistent rendering compared to 3D Gaussian splatting methods. These works focus on the general static object surface reconstruction. Within the head avatar reconstruction, NHA (Grassal et al. 2022) presents a novel approach to reconstructing full human head geometry and photorealistic textures from a short monocular RGB video. Ming et al. (Ming et al. 2024) introduces a technique for reconstructing personalized, animation-ready mesh blendshapes from single or sparse multi-view videos using neural inverse rendering. VHAP (Qian 2024) is a recent head mesh tracking tool, but it is generally used for data preprocessing tasks.
To our knowledge, we are the first to utilize Gaussians to explicitly represent head avatar meshes. Compared to 3DGS-based methods like GaussianAvatars, which exhibit poorer surface reconstruction, our approach mixes the advantages of 2DGS and 3DGS to achieve an optimal balance between rendering and reconstruction quality.
Method
Preliminaries
@@
$G(\mathbf{p})=\exp\left(-\frac{1}{2}(\mathbf{p}-\mathbf{p_{k}})^{\top}\Sigma^{-1}(\mathbf{p}-\mathbf{p_{k}})\right)$ (1)
@@
$\mathbf{c}(\mathbf{x})=\sum_{k=1}^{K}\mathbf{c_{k}}\alpha_{k}G^{3D}_{k}(\mathbf{x})\prod_{j=1}^{k-1}(1-\alpha_{j}G^{3D}_{j}(\mathbf{x}))$ (2)
@@
Overview
We propose a differentiable mixed 2D-3D Gaussian splatting method to accurately reconstruct a geometrically accurate head avatar from multi-view 2D images, as shown in Fig. 2. We begin by training 2D Gaussian representations of the 3D head avatar to ensure geometric accuracy. Next, we identify the 2D Gaussians for which the rendering quality of 2DGS is inadequate. For these identified 2D Gaussians, we bind additional 3D Gaussians to construct the mixed 2D-3D Gaussians. We use a local-to-global transformation method to convert mixed 2D-3D Gaussians for creating dynamic 3D head avatars. Finally, we train the mixed 2D-3D Gaussians to create realistic and geometrically accurate head avatars.
Mixed 2D-3D Gaussian Avatar
Mixed 2D-3D Gaussian Modeling
@@
$U=\bigcup_{i=1}^{n}S^{\prime}_{i}$ (4)
Figure 3: The tree structure of mixed 2D-3D Gaussians.
@@
1:Rendered image $\hat{X}$, Ground truth image $X$
@@
Mixed 2D-3D Gaussian Splatting
@@
$T^{2D}=\prod_{t=1}^{i-1}\left(1-\alpha_{t}G^{2D}_{t}(\mathbf{x})\right)$ (5)
$\mathbf{c}_{p}^{2D}(\mathbf{x})=\sum_{i=1}^{k}\mathbf{c_{i}}\alpha_{i}G^{2D}_{i}(\mathbf{x})T^{2D}$ (6)
$\mathbf{c}^{3D}(\mathbf{x})=\mathbf{c_{j}}\alpha_{j}G^{3D}_{j}(\mathbf{x})\prod_{t=1}^{k}\left(1-\alpha_{t}G^{2D}_{t}(\mathbf{x})\right)$ (7)
$\mathbf{c}_{b}^{2D}(\mathbf{x})=\sum_{i=k+1}^{N}\mathbf{c_{i}}\alpha_{i}G^{2D}_{i}(\mathbf{x})\left(1-\alpha_{j}G^{3D}_{j}(\mathbf{x})\right)T^{2D}$ (8)
@@
$\mathbf{c}^{mixed}(\mathbf{x})=\mathbf{c}_{p}^{2D}(\mathbf{x})+\mathbf{c}^{3D}(\mathbf{x})+\mathbf{c}_{b}^{2D}(\mathbf{x})$ (9)
Local-to-Global Transformation
To create dynamic 3D head avatars, we propose a novel local-to-global transformation method designed for mixed 2D-3D Gaussians, inspired by the 3D Gaussian rigging approach in GaussianAvatars (Qian et al. 2024). The core of the local-to-global transformation is to establish a mapping between the FLAME mesh and the mixed Gaussians. As shown in the Gaussian tree relation in Fig. 3, we associate each 2D Gaussian in local space with the triangular mesh in global space. Likewise, we connect each 3D Gaussian in local space to its corresponding 2D Gaussian. The FLAME parameters, such as expressions, enable the canonical FLAME to transform into a deformable FLAME. Through this mapping, the triangular mesh can adjust along with the mixed Gaussians, allowing these mixed Gaussians to be represented in a global space.

### v3

@@ Abstract.
Reconstructing high-fidelity 3D head avatars is crucial in various applications such as virtual reality. The pioneering methods reconstruct realistic head avatars with Neural Radiance Fields (NeRF), which have been limited by training and rendering speed. Recent methods based on 3D Gaussian Splatting (3DGS) significantly improve the efficiency of training and rendering. However, the surface inconsistency of 3DGS results in subpar geometric accuracy; later, 2DGS uses 2D surfels to enhance geometric accuracy at the expense of rendering fidelity. To leverage the benefits of both 2DGS and 3DGS, we propose a novel method named MixedGaussianAvatar for realistically and geometrically accurate head avatar reconstruction. Our main idea is to utilize 2D Gaussians to reconstruct the surface of the 3D head, ensuring geometric accuracy. We attach the 2D Gaussians to the triangular mesh of the FLAME model and connect additional 3D Gaussians to those 2D Gaussians where the rendering quality of 2DGS is inadequate, creating a mixed 2D-3D Gaussian representation. These 2D-3D Gaussians can then be animated using FLAME parameters. We further introduce a progressive training strategy that first trains the 2D Gaussians and then fine-tunes the mixed 2D-3D Gaussians. We use a unified mixed Gaussian representation to integrate the two modalities of 2D image and 3D mesh. Furthermore, the comprehensive experiments demonstrate the superiority of MixedGaussianAvatar. The code will be released.
Mixed Gaussian Splatting; Head Avatar; 2D Rendering; 3D Reconstruction
1. Introduction
@@
Reconstructing high-quality 3D head avatars from multiple 2D images captured from different viewpoints is essential for various applications (Wohlgenannt et al., 2020; Chollet et al., 2009; Bai et al., 2024; Wei et al., 2024a; Chen et al., 2025; Shao et al., 2024; Wei et al., 2025). The key high-fidelity aspects include the geometric and realistic accuracy of 3D head avatars and rendered images. With the advancement of deep learning, techniques based on Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) have gained popularity, providing significant benefits for creating high-fidelity 3D head avatars.
@@
Figure 1. MixedGaussianAvatar uses a mixed 2D-3D Gaussian Splatting method to reconstruct a realistically and geometrically accurate 3D head avatar mesh.
@@
NeRF introduces the idea of radiance fields, utilizing neural networks to store 3D information, which allows for reconstructing high-quality static scenes. When it comes to 3D head avatar reconstruction, the proposed methods primarily focus on reconstructing dynamic head avatars (Gao et al., 2022; Zielonka et al., 2022). For example, Neural Head Avatars (NHA) (Grassal et al., 2022) process RGB video inputs that display various expressions and perspectives. NHA involves estimating and refining the low-dimensional shape, expression, and pose parameters of the FLAME head model from an RGB input frame, which allows for creating a neural head avatar that can be animated and rendered from novel viewpoints. However, the head avatars generated using NeRF-based methods have lower accuracy due to the implicit representation characteristics of NeRF, and the training and rendering processes are slow.
@@
Recently, 3DGS effectively models a static scene using 3D Gaussians and rendered it through a differentiable rasterizer, greatly speeding up the 3D reconstruction process from multi-view RGB images. Because of 3DGS’s explicit representation and efficient rasterization properties, it achieves high rendering quality and fast training and inference speeds (Chen et al., 2024; Wang et al., 2025; Hu et al., 2024a, b). In 3D head avatar reconstruction, most methods attach 3D Gaussians to the FLAME triangular mesh to create dynamic facial expressions (Yu et al., 2024). For example, GaussianAvatars (Qian et al., 2024a) employ a rigging approach to align the local positions, rotations, and scales of 3D Gaussians with the global parameters. FlashAvatar (Xiang et al., 2024) and Gaussian Head Avatar  (Xu et al., 2024) utilize neural networks to predict 3D Gaussian parameters. The advantage of these methods lies in their ability to dynamically represent 3D head avatars while producing high-quality images. However, due to the inherent multi-view inconsistency of 3DGS, it cannot accurately reconstruct geometric surfaces.
@@
2D Gaussian Splatting (Huang et al., 2024b) replaces 3D Gaussians with 2D Gaussians and employs a ray-splat intersection method for the splatting process, ensuring consistency across multiple views. The advantage lies in the high quality of reconstructed geometric surfaces, but it compromises the quality of rendered images. To summarize 3DGS and 2DGS, 1) 3DGS has achieved realistic results in novel view synthesis, but it struggles to capture accurate geometric structures. 2) 2DGS uses 2D surfels to accurately reconstruct geometric surfaces, but it cannot render realistic images as effectively as 3DGS.
@@
To unify the two modalities of 2D image rendering and 3D mesh reconstruction, we harness the advantages of both 3DGS and 2DGS and introduce an innovative method called MixedGaussianAvatar. This method uses a mixed 2D-3D Gaussians representation, combining the color rendering advantages of 3DGS with the geometric reconstruction strengths of 2DGS to achieve a realistically and geometrically accurate 3D head avatar reconstruction. As shown in Fig. 1, our primary goal is to use 2DGS to preserve the geometric accuracy of the 3D head surface. We attach 2D Gaussians to the triangular mesh of the FLAME model and then connect additional 3D Gaussians to the corresponding 2D Gaussians in areas where the rendering quality is insufficient. The mixed 2D-3D Gaussians can be driven using FLAME parameters to form a dynamic 3D representation. To train the mixed 2D-3D Gaussians, we further propose a progressive training strategy that first trains the 2D Gaussians and then finetunes the mixed 2D-3D Gaussians. Our method unifies image rendering and mesh reconstruction through a mixed Gaussian representation, with contributions summarized as follows:
@@
2. Related Work
Figure 2. Pipeline of MixedGaussianAvatar. We propose a differentiable mixed 2D-3D Gaussian Splatting method for reconstructing realistically and geometrically accurate head avatars from multi-view 2D images. This approach uses 2D Gaussians for geometric precision and 3D Gaussians to correct color errors, resulting in high-quality 3D head mesh and realistic mesh textures or RGB images. The method is integrated with the FLAME model, enabling dynamic effects and animation driven by parameters through local-to-global transformation and the progressive training strategy.
@@ 2. Related Work
The representation of dynamic scenes has attracted considerable attention. Previous methods, like NeRF (Mildenhall et al., 2021) and its variants (Müller et al., 2022; Chen et al., 2022; Wei et al., 2024c, b), focus on modeling static scenes, storing them in learnable neural networks to achieve competitive novel view synthesis quality. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) and its variants (Lu et al., 2024) have introduced a much faster training and rendering pipeline. The methods mentioned above are only effective for modeling static scenes and have difficulty handling dynamic scenes. Based on NeRF and 3DGS, several approaches (Shao et al., 2023; Wu et al., 2024; Huang et al., 2024a) introduce 4D coordinates to effectively represent dynamic scenes. Humans are very sensitive to facial details, making it challenging for general methods to create high-fidelity head avatars. Recently, numerous efforts (Grassal et al., 2022; Hong et al., 2022; Chen et al., 2023) have been made to develop specialized techniques for representing 3D head avatars. AvatarMAV (Xu et al., 2023) and INSTA (Zielonka et al., 2023b), both based on voxel representations, enable fast avatar rendering. PointAvatar (Zheng et al., 2023) is the first to employ a deformable point-based representation for creating high-fidelity avatars. Inspired by 3DGS and PointAvatar, methods like GaussianAvatars (Qian et al., 2024a), Gaussian Head Avatar (Xu et al., 2024), and FlashAvatar (Xiang et al., 2024) attach 3D Gaussians to dynamic meshes or UV maps, achieving success in real-time 3D head avatar animation. Previous methods fail to consider geometric reconstruction, which makes them less appropriate for industrial applications. In contrast, MixedGaussianAvatar integrates the high-quality appearance of 3DGS with the consistent geometry of 2DGS.
@@
Compared to neural rendering, which primarily focuses on view synthesis, neural reconstruction emphasizes explicit surface extraction and textured mesh reconstruction (Aneja et al., 2025; Qiu et al., 2025; Lee et al., 2024; Moon et al., 2025; Zielonka et al., 2023a; Zhang et al., 2024; Li et al., 2024; Liu et al., 2024; Qian et al., 2024b; Metzer et al., 2023). NeuS (Wang et al., 2021) is the first to propose a bias-free volume rendering method based on neural fields that leverages a signed distance function (SDF) to achieve high-quality surface reconstruction. UNISURF (Oechsle et al., 2021) presents a unified framework that combines neural implicit surfaces and radiance fields to enable accurate 3D surface reconstruction from multi-view images without requiring object masks. VolSDF (Yariv et al., 2021) significantly improves geometry approximation and disentangles shape and appearance compared to previous neural rendering methods. With the development of 3D Gaussian Splatting, SuGaR (Guédon and Lepetit, 2024) introduces a method for fast and precise 3D mesh reconstruction and high-quality mesh rendering by aligning 3D Gaussian splats with scene surfaces, enabling efficient and detailed mesh extraction within minutes. 2D Gaussian Splatting (2DGS) (Huang et al., 2024b) proposes a novel approach for geometrically accurate radiance field reconstruction by leveraging 2D Gaussian primitives, which significantly improves surface modeling and view-consistent rendering compared to 3D Gaussian splatting methods. These works focus on the general static object surface reconstruction. Within the head avatar reconstruction, NHA (Grassal et al., 2022) presents a novel approach to reconstructing full human head geometry and photorealistic textures from a short monocular RGB video. Ming et al. (Ming et al., 2024) introduces a technique for reconstructing personalized, animation-ready mesh blendshapes from single or sparse multi-view videos using neural inverse rendering.
To our knowledge, we are the first to utilize mixed Gaussians to explicitly represent head avatar meshes. Compared to 3DGS-based methods like GaussianAvatars, which exhibit poorer surface reconstruction, our approach mixes the advantages of 2DGS and 3DGS to achieve an optimal balance between rendering and reconstruction quality.
3. Method
3.1. Preliminaries
@@
(1) $G(\mathbf{p})=\exp\left(-\frac{1}{2}(\mathbf{p}-\mathbf{p_{k}})^{\top}\Sigma^{-1}(\mathbf{p}-\mathbf{p_{k}})\right)$
@@
(2) $\mathbf{c}(\mathbf{x})=\sum_{k=1}^{K}\mathbf{c_{k}}\alpha_{k}G^{3D}_{k}(\mathbf{x})\prod_{j=1}^{k-1}(1-\alpha_{j}G^{3D}_{j}(\mathbf{x}))$
@@
3.2. Overview
We propose a differentiable mixed 2D-3D Gaussian splatting method to reconstruct a realistically and geometrically accurate head avatar from multi-view 2D images, as shown in Fig. 2. We begin by training 2D Gaussian representations of the 3D head avatar to ensure geometric accuracy. Next, we identify the 2D Gaussians for which the rendering quality of 2DGS is inadequate. For these identified 2D Gaussians, we bind additional 3D Gaussians to construct the mixed 2D-3D Gaussians in Sec. 3.3. We use a local-to-global transformation method to convert mixed 2D-3D Gaussians for creating dynamic 3D head avatars. Finally, we train the mixed 2D-3D Gaussians to create realistic and geometrically accurate head avatars. The progressive training strategy is introduced in Sec. 3.4.
3.3. Mixed 2D-3D Gaussian Avatar
3.3.1. Mixed 2D-3D Gaussian Modeling
@@
(4) $U=\bigcup_{i=1}^{n}S^{\prime}_{i}$
Figure 3. The tree structure of mixed 2D-3D Gaussians.
@@
1:Ground truth image $X$
@@
3.3.2. Mixed 2D-3D Gaussian Splatting
@@
(5) $T^{2D}=\prod_{t=1}^{i-1}\left(1-\alpha_{t}G^{2D}_{t}(\mathbf{x})\right)$
(6) $\mathbf{c}_{p}^{2D}(\mathbf{x})=\sum_{i=1}^{k}\mathbf{c_{i}}\alpha_{i}G^{2D}_{i}(\mathbf{x})T^{2D}$
(7) $\mathbf{c}^{3D}(\mathbf{x})=\mathbf{c_{j}}\alpha_{j}G^{3D}_{j}(\mathbf{x})\prod_{t=1}^{k}\left(1-\alpha_{t}G^{2D}_{t}(\mathbf{x})\right)$
(8) $\mathbf{c}_{b}^{2D}(\mathbf{x})=\sum_{i=k+1}^{N}\mathbf{c_{i}}\alpha_{i}G^{2D}_{i}(\mathbf{x})\left(1-\alpha_{j}G^{3D}_{j}(\mathbf{x})\right)T^{2D}$
@@
(9) $\mathbf{c}^{mixed}(\mathbf{x})=\mathbf{c}_{p}^{2D}(\mathbf{x})+\mathbf{c}^{3D}(\mathbf{x})+\mathbf{c}_{b}^{2D}(\mathbf{x})$
3.3.3. Local-to-Global Transformation
To create dynamic 3D head avatars, we propose a novel local-to-global transformation method designed for mixed 2D-3D Gaussians, inspired by the 3D Gaussian rigging approach in GaussianAvatars (Qian et al., 2024a). The core of the local-to-global transformation is to establish a mapping between the FLAME mesh and the mixed Gaussians. As shown in the Gaussian tree relation in Fig. 3, we associate each 2D Gaussian in local space with the triangular mesh in global space. Likewise, we connect each 3D Gaussian in local space to its corresponding 2D Gaussian. The FLAME parameters, such as expressions, enable the canonical FLAME to transform into a deformable FLAME. Through this mapping, the triangular mesh can adjust along with the mixed Gaussians, allowing these mixed Gaussians to be represented in a global space.